the sizes of computer monitors and television screens are found by measuring the length of the diagonal of the rectangular screen. the 50 inch rectangular tv below has a bottom length of 41 inches. what is the height of the screen? Moderate complexity
Answer by stanbon(75887)

the sizes of computer monitors and television screens are found by measuring the length of the diagonal of the rectangular screen. the 50 inch rectangular tv below has a bottom length of 41 inches. what is the height of the screen?
-----------------------
diag^2 = length^2 + height^2
----------------------------------
50^2 = 41^2 + height^2
-------
height = sqrt[50^2-41^2] = 28.61
====================
Cheers,
Stan H.
====================